Problem 13. Show that for any g EL (V, C) and u € V with g(u) # 0: V = null g @ {tu: 1€C}.
Answer #1

*2 Let geblr, ef &uer with giusto Then 9:V¢ is meat fonctional. we show that = nelly @ {xulxees First we prove that every ver1290-92)=0 (2) g(u)=0 12-01-0 Figuto) Aradz - xMarkell ) Vz=1 v-v=Ve-vel anv/= karto V-vao Vavi NE& V2 = h Honce every elem